# Adapting Differentially Private Synthetic Data to Relational Databases

### v1

@@ Appendix C Omitted Proofs
Below, we provide a more rigorous statement of Theorem  2 along with its proof.
Theorem 4.
@@
We run Algorithm 1 (without randomized rounding) for a single iteration across all $k$-way workloads to build $m^{\text{syn}}(\leq n_{1}^{\text{syn}}n_{2}^{\text{syn}})$ relationships for the synthetic database. Let $\bm{B}^{\text{syn}}$ be the output with privacy budget $(\varepsilon_{\text{rel}},\delta_{\text{rel}})$ and let $\bm{B}^{*}$ be the optimal bi-adjacency matrix achievable for the synthetic individual tables without any privacy constraints. We stack all the query vectors as matrices $\bm{Q}$ and $\hat{\bm{Q}}$ for the original and synthetic tables, respectively. Let $\bm{a}$ denote the true query answers computed from the original database. With probability at least $1-\beta$, we have:
@@
where $\rho_{\text{rel}}=\left(\sqrt{\varepsilon_{\text{rel}}+\operatorname{log}\frac{1}{\delta_{\text{rel}}}}-\sqrt{\operatorname{log}\frac{1}{\delta_{\text{rel}}}}\right)^{2}$.
We first prove some lemmas that will be used in the proof of Theorem 4.
Lemma 7.
(Contractive property of convex sets projection [40]) Let $\mathcal{Y}$ be a non-empty closed convex subset of $\mathbb{R}^{d}$, then for any $\bm{x}_{1},\bm{x}_{2}$ if we denote their projections by $\bm{y}_{1},\bm{y}_{2}$ respectively, i.e. $\bm{y}_{i}=\arg\min_{\bm{y}\in\mathcal{Y}}\|\bm{x}_{i}-\bm{y}\|_{2}^{2}$, then we have
@@
Proof.
Let $\bm{v}\triangleq\bm{y}_{2}-\bm{y}_{2}\neq o$. The function $f$ defined by $f(t)\triangleq|\bm{x}_{1}-(\bm{y}_{1}+t\bm{v})|^{2}$ for $t\in[0,1]$ has a minimum at $t=0$, hence $f^{\prime}(0)\geq 0$. This gives $\langle\bm{x}_{1}-\bm{y}_{1},\bm{v}\rangle\leq 0$. Similarly we obtain $\langle\bm{x}_{2}-\bm{y}_{2},\bm{v}\rangle\geq 0$. Thus, the segment $[\bm{x}_{1},\bm{x}_{2}]$ meets the two hyperplanes that are orthogonal to $\bm{v}$ and that go through $\bm{y}_{1}$ and $\bm{y}_{2}$, respectively. ∎
Lemma 8.
@@
Proof.
By the optimality of $\bm{y}^{*}$ and $\bm{y}^{\prime}\in\mathcal{Y}$, we have
@@
which yields the desired inequality. ∎
The following lemma is a standard result from concentration inequalities.
Lemma 9.
@@
Proof.
Let $u\triangleq\sqrt{2\sigma^{2}(\log n+t)}$. We have
@@
$\leq ne^{-\frac{u^{2}}{2\sigma^{2}}}=\exp(-t).$
Proof of Theorem  4.
We define $N^{\text{syn}}=n_{1}^{\text{syn}}\cdot n_{2}^{\text{syn}}$. Additionally, we define the projection set with respect to $\hat{\bm{Q}}$ as
$\mathcal{K}=\left\{\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}\ \Big{|}\ \bm{b}\in\mathbb{R}^{N^{\text{syn}}},\bm{1}^{T}\bm{b}=m^{\text{syn}},0\leq\bm{b}\leq 1\right\}.$ (11)
@@
Since $\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}^{*},\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\widetilde{\bm{b}}\in\mathcal{K}$, we have $\|\bm{b}^{*}\|_{1}=\|\widetilde{\bm{b}}\|_{1}=m^{\text{syn}}$. We denote the $j$-th column of $\hat{\bm{Q}}$ by $\bm{p}_{j}$ for $j=1,\cdots,N^{\text{syn}}$. Then
@@
$=\frac{2}{m^{\text{syn}}}\sum_{j=1}^{N^{\text{syn}}}(\widetilde{b}_{j}-b^{*}_{j})\cdot\left\langle\bm{p}_{j},\bm{w}\right\rangle$
$\leq\frac{2}{m^{\text{syn}}}\sum_{j=1}^{N^{\text{syn}}}|\widetilde{b}_{j}-b^{*}_{j}|\cdot\max_{j=1}^{N^{\text{syn}}}|\left\langle\bm{p}_{j},\bm{w}\right\rangle|$
$\leq\frac{2}{m^{\text{syn}}}\cdot(\|\widetilde{\bm{b}}\|_{1}+\|\bm{b}^{*}\|_{1})\cdot\max_{j=1}^{N^{\text{syn}}}|\left\langle\bm{p}_{j},\bm{w}\right\rangle|$
$=4\cdot\max_{j=1}^{N^{\text{syn}}}|\left\langle\bm{p}_{j},\bm{w}\right\rangle|.$
@@
$X_{j}\triangleq\langle\bm{p}_{j},\bm{w}\rangle,\ X_{N^{\text{syn}}+j}\triangleq-\langle\bm{p}_{j},\bm{w}\rangle,\quad\text{for }j=1,\cdots,N^{\text{syn}}.$
Since each random variable $X_{j}$ for $j\in[2N^{\text{syn}}]$ is a linear combination of independent zero mean Gaussian random variables, $X_{j}$ is also a zero mean Gaussian random variable. Next, we prove that $\bm{p}_{j}$ has $|\mathcal{W}_{\text{cross},k}|$ elements equal to $1$ and the remaining elements being 00. Here $|\mathcal{W}_{\text{cross},k}|$ represents the number of (cross-table) $k$-way workloads. Note that there exist $r\in[n^{\text{syn}}_{1}],s\in[n^{\text{syn}}_{2}]$ such that
@@
Recall the definition of query vector in Lemma 1. The $i$-th element of $\bm{p}_{j}$ is an indicator of whether the value of the $i$-th cross-table $k$-way marginal query of $(\bm{x}^{\text{syn}}_{r},\bm{x}^{\text{syn}}_{s})$ is $1$. Here $\bm{x}^{\text{syn}}_{r}$ and $\bm{x}^{\text{syn}}_{s}$ are the $r$-th and $s$-th records of $\mathcal{D}_{1}^{\text{syn}}$ and $\mathcal{D}_{2}^{\text{syn}}$, respectively. There are $|\mathcal{W}_{\text{cross},k}|$ workloads in total. Within each workload, for queries associated with it, only one query can have a value of $1$ for $(\bm{x}^{\text{syn}}_{r},\bm{x}^{\text{syn}}_{s})$. Hence, there are $|\mathcal{W}_{\text{cross},k}|$ elements of $\bm{p}_{j}$ being $1$. Then for $j=1,\cdots,N^{\text{syn}}$,
$\mathsf{Var}\left(X_{N^{\text{syn}}+j}\right)$ $=\mathsf{Var}\left(X_{j}\right)$
@@
$\left\|\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\widetilde{\bm{b}}-\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}^{*}\right\|_{2}^{2}$ $\leq 4\cdot\max_{j=1}^{2N^{\text{syn}}}X_{j}$
$\leq 4\cdot\sqrt{2|\mathcal{W}_{\text{cross},k}|\cdot\mathsf{Var}\left(w_{i}\right)\left(\operatorname{log}(2N^{\text{syn}})+\operatorname{log}\frac{1}{\beta}\right)}.$
@@
$\left\|\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\hat{\bm{b}}-\bm{a}\right\|_{2}\leq 2\left\|\bm{a}-\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}^{*}\right\|_{2}+\left(32|\mathcal{W}_{\text{cross},k}|\cdot\mathsf{Var}\left(w_{i}\right)\left(\operatorname{log}(2N^{\text{syn}})+\operatorname{log}\frac{1}{\beta}\right)\right)^{\frac{1}{4}}.$
Since we run Algorithm 1 (without randomized rounding) for a single iteration across all $k$-way workloads, we have $T=1$, $K=|\mathcal{W}_{\text{cross},k}|$, and $\alpha=0$. Therefore,
@@
$\left\|\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\hat{\bm{b}}-\bm{a}\right\|_{2}^{2}\leq 8\left\|\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}^{*}-\bm{a}\right\|_{2}^{2}+8\sqrt{2}|\mathcal{W}_{\text{cross},k}|\frac{d_{\text{max}}}{m\sqrt{\rho_{\text{rel}}}}\left(\operatorname{log}\frac{2N^{\text{syn}}}{\beta}\right)^{\frac{1}{2}}.$
Appendix D Discretization Strategies for Preserving Referential Integrity
Algorithm 6 can be done more efficiently based on properties of exponential random variables.
Algorithm 7 Exponential rounding for many-to-many databases.
Input: relaxed bi-adjacency matrix $\bm{B}\in[0,1]^{n_{1}^{\text{syn}}\times n_{2}^{\text{syn}}}$, number of relationship in synthetic tables: $m^{\text{syn}}$
Generate $N=n_{1}^{\text{syn}}\cdot n_{2}^{\text{syn}}$, independent random numbers, $X_{1},\cdots,X_{N}$, where $X_{i}$ is generated from $\text{Exp}(\bm{B}_{i,:})$
Sort $X_{1},\cdots,X_{N}$ from smallest to largest
Let $\mathcal{E}$ be set of indices of the $m^{\text{syn}}$ smallest numbers.
$\bm{B}=$ Reshape $\mathcal{E}^{u}$ to a $n_{1}^{\text{syn}}\times n_{2}^{\text{syn}}$ matrix
Output: unweighted bi-adjacency matrix with $m$ edges $\bm{B}$
The time complexity of this method is $O(N\log N)$ whereas the simple-and-reject approach presented in Algorithm 6 takes at least $\Omega\left(N^{2}\right)$ time.
Theorem 5.
Algorithms 6 and 7 are equivalent.
Proof.
To prove that these two procedures are equivalent, we must first remember the following properties of exponential random variables.
  1. 1.
    If $X$ is exponential with parameter $\lambda$, and $Y$ is independently exponential with parameter $\mu$, then $\min(X,Y)$ is exponential with parameter $\lambda+\mu$.
  2. 2.
    With the same assumptions as the last point, $P(X<Y)=\frac{\lambda}{\lambda+\mu}$.
  3. 3.
    The exponential distribution is memoryless, in the sense that $P(X\leq s+t\mid X>t)=P(X\leq s)$ for all $s,t\geq 0$.
This implies that, for example, the probability that the first sampled element is $X_{1}$ is
$P(X_{1}<\min(X_{2},X_{3},\ldots,X_{N}))=\frac{\hat{\bm{b}}^{r}_{1}}{\hat{\bm{b}}^{r}_{1}+\left(\hat{\bm{b}}^{r}_{2}+\cdots+\hat{\bm{b}}^{r}_{N}\right)}.$
Then, for subsequent samples, you can do a similar calculation where you leave $\hat{\bm{b}}^{r}_{1}$ out (because of memorylessness) and conclude that the probability of each potential outcome happening is similar in both algorithms. ∎
Now, we turn our attention to an adapted version of the primary algorithm tailored for databases where each entry in the first table is linked to just one entry in the second table, known as one-to-many databases. This adaptation introduces two notable distinctions from the main algorithm. Firstly, we modify the projection algorithm to seek solutions that maintain a fixed degree of one for each entry in the first dataset. We introduce a linear equation for each entry into the projection problem to accomplish this. It is worth mentioning that this adjusted projection step and the additional linear constraints remain computationally manageable and can be solved using mirror descent. Secondly, in Algorithm 8, we adopt a strategy that selects precisely one edge from the set of edges connected to each entry in the initial dataset. This approach yields an unbiased estimator for each query.
Algorithm 8 Categorical Rounding for One-to-Many databases.
Input: relaxed bi-adjacency $\bm{B}\in[0,1]^{n_{1}^{\text{syn}}\times n_{2}^{\text{syn}}}$
Assert $\operatorname{sum}(\bm{B}_{i,:})==1$ for $i\in\{1,\cdots,n_{1}^{\text{syn}}\}$
Initialize: $\bm{B}^{\text{syn}}=\bm{0}\in\mathbb{R}^{n_{1}^{\text{syn}}\times n_{2}^{\text{syn}}}$
for $i=1,\cdots,n_{1}^{\text{syn}}$ do
     Let $x_{i}$ be a random variable that takes value $j$ with probability $B_{i,j}$
     Let $B_{i,x_{i}}^{\text{syn}}=1$
end for
Output: unweighted bi-adjacency matrix $\bm{B}^{\text{syn}}$
Appendix E Extension for Improving Scalability
In this section, we introduce an extension of our main algorithm (Algorithm 1) for synthesizing large-scale relational databases. The key idea is to leverage the sparsity of the bi-adjacency matrix $\bm{B}^{\text{syn}}$—we can store $\bm{B}^{\text{syn}}$ as a sparse matrix and update a portion of its elements at each iteration. We present the extended algorithm in Algorithm 9 and illustrate its difference compared with Algorithm 1 in Figure 1.
Recall that the dimension of the bi-adjacency matrix $\bm{B}^{\text{syn}}$ is $n_{1}^{\text{syn}}\times n_{2}^{\text{syn}}$, which could result in scalability issues when generating large-scale synthetic tables. For this purpose, we observe that the nonzero entries of $\bm{B}^{\text{syn}}$ typically scale with $O(n_{1}^{\text{syn}}+n_{2}^{\text{syn}})$. For example, in the student-teacher example, each student may only be in at most $10$ classes, upper bounding the number of non-zero entries to grow linearly with the size of the students’ synthetic table.
We propose to only learn part of the bi-adjacency matrix at each iteration. We randomly select a fraction $\alpha$ of the records in each synthetic table and we slice the $\bm{B}^{\text{syn}}$ to only include the relationships to $\bm{B}^{\text{syn}}_{\text{slice}}$ between the selected records in each table. We run a similar mirror-descent algorithm to optimize the relationships between just the selected records on the queries selected in this iteration. We then re-insert these newly learned relationships into $\bm{B}^{\text{syn}}$ for the next iteration. This allows the size of the vector on which mirror descent is performed to be reduced to $\alpha^{2}n_{1}^{\text{syn}}\times n_{2}^{\text{syn}}$, saving on runtime and memory. Additionally, we observed in experiments that this approach can even produce higher fidelity results, potentially by helping mitigate some of the “over-relaxation” incurred by the mirror descent algorithm.
Figure 1: Top: diagram illustrating the original algorithm (Algorithm 1). Bottom: diagram illustrating the new algorithm (Algorithm 9).
To implement this algorithm, we begin by storing the bi-adjacency matrix as a sparse matrix $\bm{B}^{\text{syn}}\in\{0,1\}^{n_{1}^{\text{syn}}\times n_{2}^{\text{syn}}}$. We also introduce a new parameter $\alpha<1$, indicating which fraction of relations are learned in each iteration. Next, we define the slicing operation, which will be used to extract elements of $\bm{B}^{\text{syn}}$ before running the mirror-descent algorithm.
Definition 4.
Given a matrix $\bm{M}\in\mathbb{R}^{n_{1}\times n_{2}}$ and vectors $\bm{v}_{1}\in\{0,1\}^{n_{1}}$, $\bm{v}_{2}\in\{0,1\}^{n_{2}}$, we denote the slicing operation by $\bm{M}\left[\bm{v}^{1},\bm{v}^{2}\right]$. This operation yields a matrix with $\operatorname{sum}({\bm{v}^{\text{slice}}_{1}})$ rows and $\operatorname{sum}({\bm{v}^{\text{slice}}_{2}})$ columns. The $(i,j)$ entry of this matrix corresponds to the $(v^{1}_{i},v^{2}_{j})$ entry of $\bm{M}$, where $v^{1}_{i}$ is defined as the $i$-th nonzero entry of $\bm{v}^{1}$ (similarly for $v^{2}_{i}$).
The time and space complexity of the algorithm are reduced from $O(Tn_{1}^{\text{syn}}n_{2}^{\text{syn}})$ to $O(T\alpha^{2}n_{1}^{\text{syn}}n_{2}^{\text{syn}})$ by this strategy. Hence, by choosing smaller $\alpha$, we can reduce the time complexity of the algorithm.
Algorithm 9 Extension of Algorithm 1 to improve scalability.
Input: private relational database $\mathcal{B}=(\mathcal{D}_{1},\mathcal{D}_{2},\bm{B})$; privacy budget $(\varepsilon_{\text{rel}},\delta_{\text{rel}})$; individual synthetic tables $\mathcal{D}_{1}^{\text{syn}}$, $\mathcal{D}_{2}^{\text{syn}}$; workloads per iteration $K$; number of iterations $T$; number of relationship in synthetic tables: $m^{\text{syn}}$; slicing fraction $\alpha$
Initialize: $\bm{Q}=\emptyset,\hat{\bm{a}}=\emptyset$, $n_{1}^{\text{syn}}$ by $n_{2}^{\text{syn}}$ sparse matrix $\bm{B}^{\text{syn}}$ with entries $\in\{0,1\}$
for $t=1,\cdots,T$ do
     Select $K$ workloads via the exponential mechanism
     Append all queries associated with the newly selected workloads into $\bm{Q}$
     Randomly choose $\bm{v}^{\text{slice}}_{1}\in\{0,1\}^{n_{1}^{\text{syn}}}$, $\bm{v}^{\text{slice}}_{2}\in\{0,1\}^{n_{2}^{\text{syn}}}$ s.t. $\operatorname{sum}({\bm{v}^{\text{slice}}_{i}})=\operatorname{round}(\alpha n_{i}^{\text{syn}})$ for $i\in\{1,2\}$
     Set dense matrices $\bm{B}^{\text{syn}}_{\text{slice}}=\bm{B}^{\text{syn}}[\bm{v}^{\text{slice}}_{1},\bm{v}^{\text{slice}}_{2}]$ and $\bm{Q}_{\text{slice}}=\bm{Q}[\bm{v}^{\text{slice}}_{1},\bm{v}^{\text{slice}}_{2}]$
     Compute $m_{\text{slice}}=\operatorname{sum}(\bm{B}^{\text{syn}}_{\text{slice}})$
     Solve $\bm{B}^{\text{syn}}_{\text{slice}}={\arg\min}_{\begin{subarray}{c}B_{i,j}\in[0,1]\\ \operatorname{sum}(\bm{B}_{\text{slice}}^{\text{syn}})=m_{\text{slice}}\end{subarray}}\lVert\bm{Q}_{\text{slice}}\operatorname{flatten}(\bm{B}_{\text{slice}}^{\text{syn}})-\hat{\bm{a}}\rVert_{2}^{2}$ using mirror descent
     $\bm{B}^{\text{syn}}_{\text{slice}}=\operatorname{RandomizedRounding}(\bm{B}^{\text{syn}}_{\text{slice}})$
     Re-insert $\bm{B}^{\text{syn}}_{\text{slice}}$ into $\bm{B}^{\text{syn}}$: $\bm{B}^{\text{syn}}[\bm{v}^{\text{slice}}_{1},\bm{v}^{\text{slice}}_{2}]=\bm{B}^{\text{syn}}_{\text{slice}}$
end for

### v2

@@ D.7 Proof of Theorem 6
We provide a more rigorous statement of Theorem 6 along with its proof.
Theorem 8
@@
Let $\mathcal{D}^{\text{syn}}_{1}$ and $\mathcal{D}^{\text{syn}}_{2}$ be any individual synthetic tables. We stack all the $k$-way cross-table query vectors as $\hat{\bm{Q}}$. Denote by $\bm{a}$ the true answers to these queries computed from the original database. Suppose we use a privacy budget $(\epsilon_{\text{rel}},\delta_{\text{rel}})$ to run a single iteration of Algorithm 1 (without unbiased sampling), which outputs $\bm{B}^{\text{syn}}$. Let $\bm{B}^{*}$ denote the optimal weighted bi-adjacency matrix in minimizing the mean squared error when no privacy constraints are enforced. With probability at least $1-\beta$, we have
@@
where $\rho_{\text{rel}}=\left(\sqrt{\epsilon_{\text{rel}}+\operatorname{log}\frac{1}{\delta_{\text{rel}}}}-\sqrt{\operatorname{log}\frac{1}{\delta_{\text{rel}}}}\right)^{2}$ and $m$ is the number of relationships in the original database.
We introduce some useful lemmas that will be used in the proof of Theorem 8.
Lemma 14
(Contractive property of convex sets projection Schneider (2013)) Let $\mathcal{Y}$ be a non-empty closed convex subset of $\mathbb{R}^{d}$, then for any $\bm{x}_{1},\bm{x}_{2}$ if we denote their projections by $\bm{y}_{1},\bm{y}_{2}$ respectively, i.e. $\bm{y}_{i}=\arg\min_{\bm{y}\in\mathcal{Y}}\|\bm{x}_{i}-\bm{y}\|_{2}^{2}$, then we have
@@
Proof  Let $\bm{v}\triangleq\bm{y}_{2}-\bm{y}_{2}\neq o$. The function $f$ defined by $f(t)\triangleq|\bm{x}_{1}-(\bm{y}_{1}+t\bm{v})|^{2}$ for $t\in[0,1]$ has a minimum at $t=0$, hence $f^{\prime}(0)\geq 0$. This gives $\langle\bm{x}_{1}-\bm{y}_{1},\bm{v}\rangle\leq 0$. Similarly we obtain $\langle\bm{x}_{2}-\bm{y}_{2},\bm{v}\rangle\geq 0$. Thus, the segment $[\bm{x}_{1},\bm{x}_{2}]$ meets the two hyperplanes that are orthogonal to $\bm{v}$ and that go through $\bm{y}_{1}$ and $\bm{y}_{2}$, respectively.
Lemma 15
@@
Proof  By the optimality of $\bm{y}^{*}$ and $\bm{y}^{\prime}\in\mathcal{Y}$, we have
@@
which yields the desired inequality.
The following lemma is a standard result from concentration inequalities (Boucheron et al., 2003).
Lemma 16
@@
Proof  Let $u\triangleq\sqrt{2\sigma^{2}(\log n+t)}$. We have
@@
$\leq n\exp\left(-\frac{u^{2}}{2\sigma^{2}}\right)=\exp(-t),$
which yields the desired inequality.
Next, we provide the proof of Theorem  8
Proof  We define $N=n_{1}^{\text{syn}}\cdot n_{2}^{\text{syn}}$. Additionally, we define the projection set with respect to $\hat{\bm{Q}}$ as
$\mathcal{K}=\left\{\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}\ \Big{|}\ \bm{b}\in\mathbb{R}^{N},\bm{1}^{T}\bm{b}=m^{\text{syn}},0\leq\bm{b}\leq 1\right\}.$ (26)
@@
Since $\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}^{*},\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\widetilde{\bm{b}}\in\mathcal{K}$, we have $\|\bm{b}^{*}\|_{1}=\|\widetilde{\bm{b}}\|_{1}=m^{\text{syn}}$. We denote the $j$-th column of $\hat{\bm{Q}}$ by $\bm{p}_{j}$ for $j=1,\cdots,N$. Then
@@
$=\frac{2}{m^{\text{syn}}}\sum_{j=1}^{N}(\widetilde{b}_{j}-b^{*}_{j})\cdot\left\langle\bm{p}_{j},\bm{w}\right\rangle$
$\leq\frac{2}{m^{\text{syn}}}\sum_{j=1}^{N}|\widetilde{b}_{j}-b^{*}_{j}|\cdot\max_{j=1}^{N}|\left\langle\bm{p}_{j},\bm{w}\right\rangle|$
$\leq\frac{2}{m^{\text{syn}}}\cdot(\|\widetilde{\bm{b}}\|_{1}+\|\bm{b}^{*}\|_{1})\cdot\max_{j=1}^{N}|\left\langle\bm{p}_{j},\bm{w}\right\rangle|$
$=4\cdot\max_{j=1}^{N}|\left\langle\bm{p}_{j},\bm{w}\right\rangle|.$
@@
$X_{j}\triangleq\langle\bm{p}_{j},\bm{w}\rangle,\ X_{N+j}\triangleq-\langle\bm{p}_{j},\bm{w}\rangle,\quad\text{for }j=1,\cdots,N.$
Since each random variable $X_{j}$ for $j\in[2N]$ is a linear combination of independent zero mean Gaussian random variables, $X_{j}$ is also a zero mean Gaussian random variable. Next, we prove that $\bm{p}_{j}$ has $|\mathcal{W}_{\text{cross},k}|$ elements equal to $1$ and the remaining elements being 00. Here $|\mathcal{W}_{\text{cross},k}|$ represents the number of (cross-table) $k$-way workloads. Note that there exist $r\in[n^{\text{syn}}_{1}],s\in[n^{\text{syn}}_{2}]$ such that
@@
Recall the definition of query vector in Lemma 1. The $i$-th element of $\bm{p}_{j}$ is an indicator of whether the value of the $i$-th cross-table $k$-way marginal query of $(\bm{x}^{\text{syn}}_{r},\bm{x}^{\text{syn}}_{s})$ is $1$. Here $\bm{x}^{\text{syn}}_{r}$ and $\bm{x}^{\text{syn}}_{s}$ are the $r$-th and $s$-th records of $\mathcal{D}_{1}^{\text{syn}}$ and $\mathcal{D}_{2}^{\text{syn}}$, respectively. There are $|\mathcal{W}_{\text{cross},k}|$ workloads in total. Within each workload, for queries associated with it, only one query can have a value of $1$ for $(\bm{x}^{\text{syn}}_{r},\bm{x}^{\text{syn}}_{s})$. Hence, there are $|\mathcal{W}_{\text{cross},k}|$ elements of $\bm{p}_{j}$ being $1$. Then for $j=1,\cdots,N$,
$\mathsf{Var}\left(X_{N+j}\right)$ $=\mathsf{Var}\left(X_{j}\right)$
@@
$\left\|\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\widetilde{\bm{b}}-\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}^{*}\right\|_{2}^{2}$ $\leq 4\cdot\max_{j=1}^{2N}X_{j}$
$\leq 4\cdot\sqrt{2|\mathcal{W}_{\text{cross},k}|\cdot\mathsf{Var}\left(w_{i}\right)\left(\operatorname{log}(2N)+\operatorname{log}\frac{1}{\beta}\right)}.$
@@
$\left\|\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\hat{\bm{b}}-\bm{a}\right\|_{2}\leq 2\left\|\bm{a}-\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}^{*}\right\|_{2}+\left(32|\mathcal{W}_{\text{cross},k}|\cdot\mathsf{Var}\left(w_{i}\right)\left(\operatorname{log}(2N)+\operatorname{log}\frac{1}{\beta}\right)\right)^{\frac{1}{4}}.$
Since we run Algorithm 1 (without unbiased sampling) for a single iteration across all $k$-way workloads, we have $T=1$, $K=|\mathcal{W}_{\text{cross},k}|$, and $\alpha=0$. Therefore,
@@
$\left\|\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\hat{\bm{b}}-\bm{a}\right\|_{2}^{2}\leq 8\left\|\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}^{*}-\bm{a}\right\|_{2}^{2}+8\sqrt{2}|\mathcal{W}_{\text{cross},k}|\frac{d_{\text{max}}}{m\sqrt{\rho_{\text{rel}}}}\left(\operatorname{log}\frac{2N}{\beta}\right)^{\frac{1}{2}}.$
Acknowledgments
This work was supported in part by the MIT-IBM Watson AI Lab. The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing computing resources that have contributed to the results reported within this paper.